OpenStudy (anonymous):

Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. vertex at (–3, 0) and co-vertex at (0, 2)

terenzreignz (terenzreignz):

Hi :) Whenever you have an ellipse centred around the origin (0 , 0) and with a vertex (or co-vertex) at (a , 0) and a co-vertex (or vertex) at (0 , b) the equation of your ellipse is given by this formula... \[\huge \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]